The following selected information is available for two competitors, Nike, Inc. and Adidas AG: Adidas (in euros) Nike in millions, except for per-share information) (in US.) Market price per share Dividends declared and paid Dividends per share Net income $62.50 89.91 320 1.60 634 931 1.08 3,273 Calculate the (1) payout and (2) dividend yield ratios for each company. (Round answers to 1 decimal place, eg 52%) Nike Adidas (1) Payout ratio (2) Dividend yield Which company would investors favour for dividend income purposes? Investors would likely prefer for dividend income purposes.
Answer #1

1. Payout ratio = (Dividend paid / Net Income) x 100
Nike = (931 / 3273) x 100 = 28.4%
Adidas = (320 / 634) x 100 = 50.5%

2. Dividend yield = Dividend per share / Market price per share
Nike = (1.08 / 62.50) x 100 = 1.7%
Adidas = (1.60 / 89.91) x 100 = 1.8%

Investors would likely prefer Adidas for dividend income purposes.